# Learning with Noisy Labels: the Exploration of Error Bounds in Classification

### v1

@@ 5.1 Noise-tolerance
$\mathcal{L}^{\eta}_{n}(f)$ $=\frac{1}{n}\sum^{n}_{i=1}\ell(f(x_{i}),y^{\eta}_{i})=\frac{1}{n}\sum^{n}_{i=1}\left[(1-\eta)\ell(f(x_{i}),y_{i})+\frac{\eta}{K-1}\sum^{K}_{j=1}\sum_{y_{i}\neq e_{j}}\ell(f(x_{i}),e_{j})\right]$
$=(1-\eta)\mathcal{L}_{n}(f)+\frac{\eta}{K-1}(C_{0}-\mathcal{L}_{n}(f))=\frac{C_{0}\eta}{K-1}+\left(1-\frac{\eta K}{K-1}\right)\mathcal{L}_{n}(f),$
where the third equality is from Equation (5.1). Since $\eta<\frac{K-1}{K}$ and $\hat{f_{n}}$ is the empirical minimizer for the true data in (3.3), then we have
$\mathcal{L}^{\eta}_{n}(\hat{f}_{n})-\mathcal{L}^{\eta}_{n}(f)=\left(1-\frac{\eta K}{K-1}\right)(\mathcal{L}_{n}(\hat{f}_{n})-\mathcal{L}_{n}(f))\leq 0,\ \forall f\in{\mathcal{F}},$
For the loss functions
$l(\hat{y},y)=\left\{\begin{array}[]{ll}\|\hat{y}-y\|_{1},&\ell_{1}\text{ loss},\\ -\sum_{i=1}^{k}y_{i}\log(\hat{y_{i}}),&\text{CE loss},\\ -\sum_{i=1}^{k}\hat{y_{i}}\log(y_{i}),&\text{RCE loss},\end{array}\right.$
we have the following remarks:
Remark 5.2.
Both the $\ell_{1}$ loss and the RCE loss are symmetric loss functions when we choose $f$ as a function from ${\mathbb{R}}^{d}$ to $\mathcal{Y}=\{y=[y_{1}\ \cdots\ y_{K}]^{\top}|y_{i}\geq 0,\sum^{K}_{i=1}=1\}$ [GKS17, WMC+19]. For CE loss, it is also a symmetric loss function when we set $f$ to be a function mapping ${\mathbb{R}}^{d}$ to =$\mathcal{E}=\{e_{1},\cdots,e_{K}\}$.
5.2 Error analysis
In this subsection, we analyze the error bounds for different types of loss functions, including the $\ell_{1}$ loss, the CE and RCE losses. We start from the Lipschitz-type property for the CE and RCE losses.
@@ 5.2 Error analysis
When Assumption 4.1 holds, both the CE loss function and the RCE loss function satisfy the following property
$\left|\ell(a,q)-\ell(b,q)\right|\leq\lambda\|a-b\|_{1},\ \forall a,b,q\in\mathcal{Y},$
where
$\lambda=\left\{\begin{array}[]{ll}\frac{1}{\epsilon},&\mathrm{CE\ loss},\\ -A,&\mathrm{RCE\ loss}.\end{array}\right.$ (5.2)
@@
For CE loss function $\ell(a,q)=-\sum_{i=1}^{K}q_{i}\log a_{i}$, $\ell(b,q)=-\sum_{i=1}^{K}q_{i}\log b_{i}$, we have
$\left|\ell(a,q)-\ell(b,q)\right|$ $=\left|\sum_{i=1}^{K}q_{i}\log a_{i}-\sum_{i=1}^{K}q_{i}\log b_{i}\right|=\left|\sum_{i=1}^{K}q_{i}[\log a_{i}-\log b_{i}]\right|$
$\leq\sum_{i=1}^{K}\left|q_{i}\right|\left|\log a_{i}-\log b_{i}\right|=\sum_{i=1}^{K}\left|q_{i}\right|(\log x)^{\prime}|_{x=\xi_{i}}\left|a_{i}-b_{i}\right|,(\exists\xi_{i}\in[a_{i},b_{i}])$
$\leq\frac{1}{\epsilon}\|q\|_{2}\|a-b\|_{2}\leq\frac{1}{\epsilon}\|a-b\|_{1}.$
where the third equality is from the Lagrange mean value theorem and $\max_{x}(\log x)^{\prime}=1/\epsilon$ when Assumption 4.1 holds.
For RCE loss function $\ell(a,q)=-\sum_{i=1}^{K}a_{i}\log q_{i}$, $\ell(b,q)=-\sum_{i=1}^{K}b_{i}\log q_{i}$, we directly calculate :
$\left|\ell(a,q)-\ell(b,q)\right|$ $=\left|\sum_{i=1}^{K}a_{i}\log q_{i}-\sum_{i=1}^{K}b_{i}\log q_{i}\right|=\left|\sum_{i=1}^{K}\log q_{i}(a_{i}-b_{i})\right|$
$\leq\sum_{i=1}^{K}\left|\log q_{i}\right|\left|(a_{i}-b_{i})\right|\leq\max_{i}|\log q_{i}|\|a-b\|_{1}\leq-A\|a-b\|_{1},$
where we use the fact that $\max_{i}|\log q_{i}|\leq-A$ based on Assumption 4.1. ∎
Lemma 5.2.
Let $f_{0}:[0,1]^{d}\to\mathcal{Y}=\{y=[y_{1}\ \cdots\ y_{K}]^{\top}|y_{i}\geq 0,\sum^{K}_{i=1}=1\}$ represent the true classifier, and $\hat{f_{n}}$ denote the empirical risk minimizer on the true data. Let $S=\{(x_{1},y_{1}),\cdots,(x_{n},y_{n})\in X\times Y\}$. We have the following
(1) When the loss function is the $\ell_{1}$ loss
$|\mathcal{L}(\hat{f_{n}})-\mathcal{L}(f_{0})|\leq 2\mathbb{E}_{S}\left[\sup_{f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})}\left|\mathcal{L}(f)-\mathcal{L}_{n}(f)\right|\right]+\inf_{f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})}\left\|f-f_{0}\right\|_{L^{1}([0,1]^{d})}.$ (5.3)
(2) When Assumption 4.1 holds and the loss function is either the CE loss or the RCE loss
$|\mathcal{L}(\hat{f}_{n})-\mathcal{L}(f_{0})|\leq 2\mathbb{E}_{S}\left[\sup_{f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})}\left|\mathcal{L}(f)-\mathcal{L}_{n}(f)\right|\right]+\lambda\inf_{f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})}\left\|f-f_{0}\right\|_{L^{1}([0,1]^{d})},$ (5.4)
where $\lambda$ is defined in (5.2).
Proof.
@@
Let $\tilde{f}\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})$ satisfying $\|\tilde{f}-f_{0}\|_{L^{1}([0,1]^{d})}\leq\|f-f_{0}\|_{L^{1}([0,1]^{d})},\ \forall f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})$, then
$|\mathcal{L}(\hat{f}_{n})-\mathcal{L}(f_{0})|$ $=\mathbb{E}_{S}(|\mathcal{L}(\hat{f}_{n})-\mathcal{L}(f_{0})|)$ (5.5)
$=\mathbb{E}_{S}(|\mathcal{L}(\hat{f}_{n})-\mathcal{L}_{n}(\hat{f}_{n})+\mathcal{L}_{n}(\hat{f}_{n})-\mathcal{L}_{n}(\tilde{f})+\mathcal{L}_{n}(\tilde{f})-\mathcal{L}(\tilde{f})+\mathcal{L}(\tilde{f})-\mathcal{L}(f_{0})|)$
@@
$\leq\mathbb{E}_{S}\left|\mathcal{L}(\hat{f}_{n})-\mathcal{L}_{n}(\hat{f}_{n})\right|+\mathbb{E}_{S}\left|\mathcal{L}_{n}(\tilde{f})-\mathcal{L}(\tilde{f})\right|+\left|\mathcal{L}(\tilde{f})-\mathcal{L}(f_{0})\right|$
$\leq 2\mathbb{E}_{S}\left[\sup_{f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})}\left|\mathcal{L}(f)-\mathcal{L}_{n}(f)\right|\right]+\left|\mathcal{L}(\tilde{f})-\mathcal{L}(f_{0})\right|.$
@@
Here, we utilize the fact that $\mathcal{L}_{n}(\hat{f}_{n})\leq\mathcal{L}_{n}(f),\ \forall f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})$ in the first inequality of (5.5). For the $\ell_{1}$ loss function, we have
$|\mathcal{L}(\tilde{f})-\mathcal{L}(f_{0})|$ $=|\mathbb{E}_{X,Y}\|\tilde{f}(X)-Y\|_{1}-\mathbb{E}_{X,Y}\|{f_{0}}(X)-Y\|_{1}|\leq\mathbb{E}_{X,Y}[|\|\tilde{f}(X)-Y\|_{1}-\|{f_{0}}(X)-Y\|_{1}|]$ (5.6)
$\leq\mathbb{E}_{X}\mathbb{E}_{Y|X}[\|\tilde{f}(X)-f_{0}(X)\|_{1}]=\mathbb{E}_{X}\|\tilde{f}(X)-f_{0}(X)\|_{1}=\left\|\tilde{f}-f_{0}\right\|_{L^{1}([0,1]^{d})}$
$=\inf_{f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})}\left\|f-f_{0}\right\|_{L^{1}([0,1]^{d})},$
where the second inequality is from the triangle inequality. For the CE loss function or the RCE loss function, we have
$|\mathcal{L}(\tilde{f})-\mathcal{L}(f_{0})|$ $=\left|\mathbb{E}_{X,Y}\left[\ell(\tilde{f}(X),Y)\right]-\mathbb{E}_{X,Y}\left[\ell(f_{0}(X),Y)\right]\right|\leq\lambda\mathbb{E}_{X}\|\tilde{f}(X)-f_{0}(X)\|_{1}$ (5.7)
$\leq\lambda\left\|\tilde{f}-f_{0}\right\|_{L^{1}([0,1]^{d})}=\lambda\inf_{f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})}\left\|f-f_{0}\right\|_{L^{1}([0,1]^{d})},$
where the second inequality is from Lemma 5.1 and $\lambda$ is defined in (5.2). Combining (5.5), (5.6) and (5.7), we conclude the result. ∎
For the excess risk about the empirical risk minimizer on the noisy distribution $\mathcal{L}^{\eta}(\hat{f^{\eta}_{n}})-\mathcal{L}^{\eta}(f_{0})$, we have the following similar results.
Lemma 5.3.
Let $f_{0}$ represent the true classifier, and $\hat{f^{\eta}_{n}}$ denote the empirical risk minimizer on the noisy data. Let $S^{\eta}=\{(x_{1},y^{\eta}_{1}),\cdots,(x_{n},y^{\eta}_{n})\in X\times Y^{\eta}\}$. We have the following
(1) When the loss function is the $\ell_{1}$ loss
$|\mathcal{L}^{\eta}(\hat{f^{\eta}_{n}})-\mathcal{L}^{\eta}(f_{0})|\leq 2\mathbb{E}_{S^{\eta}}\left[\sup_{f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})}\left|\mathcal{L}^{\eta}(f)-\mathcal{L}^{\eta}_{n}(f)\right|\right]+\inf_{f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})}\left\|f-f_{0}\right\|_{L^{1}([0,1]^{d})}.$ (5.8)
(2) When Assumption 4.1 holds and the loss function is either the CE los or the RCE loss
$|\mathcal{L}^{\eta}(\hat{f^{\eta}_{n}})-\mathcal{L}^{\eta}(f_{0})|\leq 2\mathbb{E}_{S^{\eta}}\left[\sup_{f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})}\left|\mathcal{L}^{\eta}(f)-\mathcal{L}^{\eta}_{n}(f)\right|\right]+\lambda\inf_{f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})}\left\|f-f_{0}\right\|_{L^{1}([0,1]^{d})},$ (5.9)
where $\lambda$ is defined in (5.2).
Proof.
Let $\tilde{f}\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})$ satisfying $\|\tilde{f}-f_{0}\|_{L^{1}([0,1]^{d})}\leq\|f-f_{0}\|_{L^{1}([0,1]^{d})},\ \forall f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})$, then
$|\mathcal{L}^{\eta}(\hat{f^{\eta}_{n}})-\mathcal{L}^{\eta}(f_{0})|$ $=\mathbb{E}_{S^{\eta}}(|\mathcal{L}^{\eta}(\hat{f^{\eta}_{n}})-\mathcal{L}^{\eta}(f_{0})|)$ (5.10)
$=\mathbb{E}_{S^{\eta}}(|\mathcal{L}^{\eta}(\hat{f^{\eta}_{n}})-\mathcal{L}^{\eta}_{n}(\hat{f^{\eta}_{n}})+\mathcal{L}^{\eta}_{n}(\hat{f^{\eta}_{n}})-\mathcal{L}^{\eta}_{n}(\tilde{f})+\mathcal{L}^{\eta}_{n}(\tilde{f})-\mathcal{L}^{\eta}(\tilde{f})+\mathcal{L}^{\eta}(\tilde{f})-\mathcal{L}^{\eta}(f_{0})|)$
$\leq\mathbb{E}_{S^{\eta}}\left|\mathcal{L}^{\eta}(\hat{f^{\eta}_{n}})-\mathcal{L}^{\eta}_{n}(\hat{f^{\eta}_{n}})\right|+\mathbb{E}_{S}\left|\mathcal{L}^{\eta}_{n}(\tilde{f})-\mathcal{L}^{\eta}(\tilde{f})\right|+\left|\mathcal{L}^{\eta}(\tilde{f})-\mathcal{L}^{\eta}(f_{0})\right|$
$\leq 2\mathbb{E}_{S^{\eta}}\left[\sup_{f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})}\left|\mathcal{L}^{\eta}(f)-\mathcal{L}^{\eta}_{n}(f)\right|\right]+\left|\mathcal{L}^{\eta}(\tilde{f})-\mathcal{L}^{\eta}(f_{0})\right|.$
Here, we utilize the fact that $\mathcal{L}^{\eta}_{n}(\hat{f^{\eta}_{n}})\leq\mathcal{L}^{\eta}_{n}(f),\ \forall f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})$ in the first inequality of (5.10). For the $\ell_{1}$ loss function, we have
$|\mathcal{L}^{\eta}(\tilde{f})-\mathcal{L}^{\eta}(f_{0})|$ $=|\mathbb{E}_{X,Y^{\eta}}\|\tilde{f}(X)-Y^{\eta}\|_{1}-\mathbb{E}_{X,Y^{\eta}}\|{f_{0}}(X)-Y^{\eta}\|_{1}|$ (5.11)
$\leq\mathbb{E}_{X,Y^{\eta}}[|\|\tilde{f}(X)-Y^{\eta}\|_{1}-\|{f_{0}}(X)-Y^{\eta}\|_{1}|]$
$\leq\mathbb{E}_{X}\mathbb{E}_{Y^{\eta}|X}[\|\tilde{f}(X)-f_{0}(X)\|_{1}]=\mathbb{E}_{X}\|\tilde{f}(X)-f_{0}(X)\|_{1}=\left\|\tilde{f}-f_{0}\right\|_{L^{1}([0,1]^{d})}$
$=\inf_{f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})}\left\|f-f_{0}\right\|_{L^{1}([0,1]^{d})},$
where the second inequality is from the triangle inequality. For the CE loss function or the RCE loss function, we have
$|\mathcal{L}^{\eta}(\tilde{f})-\mathcal{L}^{\eta}(f_{0})|$ $=\mathbb{E}_{X,Y^{\eta}}\left[\ell(\tilde{f}(X),Y^{\eta})\right]-\mathbb{E}_{X,Y^{\eta}}\left[\ell(f_{0}(X),Y^{\eta})\right]\leq\lambda\mathbb{E}_{X}\|\tilde{f}(X)-f_{0}(X)\|_{1}$ (5.12)
$\leq\lambda\left\|\tilde{f}-f_{0}\right\|_{L^{1}([0,1]^{d})}=\lambda\inf_{f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})}\left\|f-f_{0}\right\|_{L^{1}([0,1]^{d})},$
where the second inequality is from Lemma 5.1 and $\lambda$ is defined in (5.2). Combining (5.10), (5.11) and (5.12), we conclude the result. ∎
@@
We denote $\mathbb{E}_{S}\left[\sup_{f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})}\left|\mathcal{L}^{\eta}(f)-\mathcal{L}^{\eta}_{n}(f)\right|\right]$ or $\mathbb{E}_{S^{\eta}}\left[\sup_{f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})}\left|\mathcal{L}^{\eta}(f)-\mathcal{L}^{\eta}_{n}(f)\right|\right]$ as the statistical error and $\left\|f-f_{0}\right\|_{L^{1}([0,1]^{d})}$ as the approximation error. In the following subsections, we will explore both types of errors.

### v2

@@ 5.1 Error analysis
Let $\tilde{f}=\mathrm{softmax}\circ\tilde{\kappa}$ with $\tilde{\kappa}\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D},{\mathcal{M}})$ satisfying $\|\tilde{\kappa}-\kappa_{0}\|_{L^{2}([0,1]^{d})}\leq\|\kappa-\kappa_{0}\|_{L^{2}([0,1]^{d})},\ \forall\kappa\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D},{\mathcal{M}})$, then
$|\mathcal{L}^{\eta}(\hat{f^{\eta}_{n}})-\mathcal{L}^{\eta}(f_{0})|$ (5.2)
$=$ $|\mathcal{L}^{\eta}(\hat{f^{\eta}_{n}})-\mathcal{L}^{\eta}_{n}(\hat{f^{\eta}_{n}})+\mathcal{L}^{\eta}_{n}(\hat{f^{\eta}_{n}})-\mathcal{L}^{\eta}_{n}(\tilde{f})+\mathcal{L}^{\eta}_{n}(\tilde{f})-\mathcal{L}^{\eta}(\tilde{f})+\mathcal{L}^{\eta}(\tilde{f})-\mathcal{L}^{\eta}(f_{0})|$
$\leq$ $\left|\mathcal{L}^{\eta}(\hat{f^{\eta}_{n}})-\mathcal{L}^{\eta}_{n}(\hat{f^{\eta}_{n}})\right|+\left|\mathcal{L}^{\eta}_{n}(\tilde{f})-\mathcal{L}^{\eta}(\tilde{f})\right|+\left|\mathcal{L}^{\eta}(\tilde{f})-\mathcal{L}^{\eta}(f_{0})\right|$
$\leq$ $2\left[\sup_{\kappa\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D},{\mathcal{M}})}\left|\mathcal{L}^{\eta}(\mathrm{softmax}\circ\kappa)-\mathcal{L}^{\eta}_{n}(\mathrm{softmax}\circ\kappa)\right|\right]+\left|\mathcal{L}^{\eta}(\tilde{f})-\mathcal{L}^{\eta}(f_{0})\right|.$
Here, we utilize the fact that $\mathcal{L}^{\eta}_{n}(\hat{f^{\eta}_{n}})\leq\mathcal{L}^{\eta}_{n}(f),\ \forall f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D},{\mathcal{M}})$ in the first inequality of (5.2). We have
$|\mathcal{L}^{\eta}(\tilde{f})-\mathcal{L}^{\eta}(f_{0})|$ (5.3)
$=$ $\mathbb{E}_{X,Y^{\eta}}\left[\ell(\tilde{f}(X),Y^{\eta})\right]-\mathbb{E}_{X,Y^{\eta}}\left[\ell(f_{0}(X),Y^{\eta})\right]\leq\lambda\mathbb{E}_{X}\|\tilde{\kappa}(X)-\kappa_{0}(X)\|_{2}$
$\leq$ $\lambda\left\|\tilde{\kappa}-\kappa_{0}\right\|_{L^{2}([0,1]^{d})}=\lambda\inf_{\kappa\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D},{\mathcal{M}})}\left\|\kappa-\kappa_{0}\right\|_{L^{2}([0,1]^{d})}.$
Combining (5.2), and (LABEL:C.3), we conclude the result. ∎
For the excess risk about the empirical risk minimizer on the true data $\mathcal{L}(\hat{f^{\eta}_{n}})-\mathcal{L}(f_{0})$, we have the following similar results.
@@
Let $f_{0}=\mathrm{softmax}\circ\kappa_{0}:[0,1]^{d}\to\mathcal{Y}=\{y=[y_{1}\ \cdots\ y_{K}]^{\top}|y_{i}\geq 0,\sum^{K}_{i=1}=1\}$ represent the true classifier, and $\hat{f_{n}}$ denote the empirical risk minimizer on the true data. Let $|\ell(\mathrm{softmax}(a),q)-\ell(\mathrm{softmax}(b),q)|\leq\lambda\|a-b\|_{2},\ \forall a,b\in{\mathbb{R}}^{K},\forall q\in\mathcal{Y}$, the excess risks for the empirical risk minimizer $\hat{f_{n}}=\mathrm{softmax}\circ\hat{\kappa}_{n}$ with $\hat{\kappa}_{n}\in{\mathcal{F}}_{d,K}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})$. We have the following
$|\mathcal{L}(\hat{f}_{n})-\mathcal{L}(f_{0})|\leq 2\sup_{\begin{subarray}{c}f=\mathrm{softmax}\circ\kappa,\\ \kappa\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D},{\mathcal{M}})\end{subarray}}\left|\mathcal{L}(f)-\mathcal{L}_{n}(f)\right|+\lambda\inf_{\kappa\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D},{\mathcal{M}})}\left\|\kappa-\kappa_{0}\right\|_{L^{2}([0,1]^{d})}.$ (5.4)
@@
Let $\tilde{f}=\mathrm{softmax}\circ\tilde{\kappa}$ with $\tilde{\kappa}\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D},{\mathcal{M}})$ satisfying $\|\tilde{\kappa}-\kappa_{0}\|_{L^{2}([0,1]^{d})}\leq\|\kappa-\kappa_{0}\|_{L^{2}([0,1]^{d})},\ \forall\kappa\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D},{\mathcal{M}})$, then
$|\mathcal{L}(\hat{f}_{n})-\mathcal{L}(f_{0})|$ (5.5)
$=$ $|\mathcal{L}(\hat{f}_{n})-\mathcal{L}_{n}(\hat{f}_{n})+\mathcal{L}_{n}(\hat{f}_{n})-\mathcal{L}_{n}(\tilde{f})+\mathcal{L}_{n}(\tilde{f})-\mathcal{L}(\tilde{f})+\mathcal{L}(\tilde{f})-\mathcal{L}(f_{0})|$
$\leq$ $\left|\mathcal{L}(\hat{f}_{n})-\mathcal{L}_{n}(\hat{f}_{n})\right|+\left|\mathcal{L}_{n}(\tilde{f})-\mathcal{L}(\tilde{f})\right|+\left|\mathcal{L}(\tilde{f})-\mathcal{L}(f_{0})\right|$
@@
$\leq$ $2\left[\sup_{\kappa\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D},{\mathcal{M}})}\left|\mathcal{L}(\mathrm{softmax}\circ\kappa)-\mathcal{L}_{n}(\mathrm{softmax}\circ\kappa)\right|\right]+\left|\mathcal{L}(\tilde{f})-\mathcal{L}(f_{0})\right|.$
@@
Here, we utilize the fact that $\mathcal{L}_{n}(\hat{f}_{n})\leq\mathcal{L}_{n}(f),\ \forall f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D},{\mathcal{M}})$ in the first inequality of (LABEL:eq:error_decomposition). Then
$|\mathcal{L}(\tilde{f})-\mathcal{L}(f_{0})|$ $=\left|\mathbb{E}_{X,Y}\left[\ell(\tilde{f}(X),Y)\right]-\mathbb{E}_{X,Y}\left[\ell(f_{0}(X),Y)\right]\right|\leq\lambda\mathbb{E}_{X}\|\tilde{\kappa}(X)-\kappa_{0}(X)\|_{2}$ (5.6)
$\leq\lambda\left\|\tilde{\kappa}-\kappa_{0}\right\|_{L^{2}([0,1]^{d})}=\lambda\inf_{\kappa\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D},{\mathcal{M}})}\left\|\kappa-\kappa_{0}\right\|_{L^{2}([0,1]^{d})}.$
Combining (LABEL:eq:error_decomposition), and (5.6), we conclude the result. ∎
@@
We denote $\sup_{f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D},{\mathcal{M}})}\left|\mathcal{L}^{\eta}(f)-\mathcal{L}^{\eta}_{n}(f)\right|$ or $\sup_{f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D},{\mathcal{M}})}\left|\mathcal{L}(f)-\mathcal{L}_{n}(f)\right|$ as the statistical error and $\inf_{\kappa\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D},{\mathcal{M}})}\left\|\kappa-\kappa_{0}\right\|_{L^{2}([0,1]^{d})}$ as the approximation error. In the following subsections, we will explore both types of errors.